OpenStudy (anonymous):

solve x^2=2x=15 by completing the square a. 3,-5 b. 6,-2 c. 5, -15 d. 4, -15

OpenStudy (kewlgeek555):

Okay, so if you are a fast math problem solver, try to plug the values in. I'll dot he first one for you:\[x ^{2}=2x=15\]\[3^{2}=2\times-5=15\]\[9=2\times-5=15\]Since you already see that one of them is 9, then you know that a is not the answer. Try the other ones.

OpenStudy (anonymous):

so i plug in one of them to the equation?! @kewlgeek555

OpenStudy (kewlgeek555):

Yes. You plug in B, just like I

OpenStudy (kewlgeek555):

did with A.

OpenStudy (anonymous):

on the beginning part why is x^2 =2x=15 <<< where that from

OpenStudy (kewlgeek555):

The directions say that the two first equations have to equal 15, so when you plug in the values, they have to be 15 = 15 = 15.

OpenStudy (kewlgeek555):

Here - let me help you do B: \[x ^{2}=2x=15\]\[6^{2}=2\times-2=15\]\[36=2\times-2=15?\]FALSE. The answer is NOT B or A.

OpenStudy (kewlgeek555):

Try to do C by yourself.

OpenStudy (anonymous):

omg!! lol its supposed to be x^2+2x=15!! my bad